OpenStudy (omgthatgirl):

Which of the following is the correct notation for the complex number sqrt-81-14 A. -14+9i B. -14+sqrt-81 C. -14+i sqrt 9 D. 14-9i

OpenStudy (omgthatgirl):

Yes thats right

OpenStudy (jhannybean):

\[\sqrt{-81}-14\]\[\sqrt{-1}\sqrt{81}-14\]\[i=\sqrt{-1}\]\[i\sqrt{81}-14\]

OpenStudy (jhannybean):

And then reduce what \(\sqrt{81}\) is.

OpenStudy (omgthatgirl):

9

OpenStudy (omgthatgirl):

@Jhannybean

OpenStudy (omgthatgirl):

Socwould it be sqrt 9